"Ms. Kremer would like to purchase a new condo for $102,000. She plans to make a down payment of $55,000 and to borrow the rest of the money from the bank. The bank charges a nominal annual interest rate of 4% compounded daily. She agrees to monthly payments to pay off the loan in 12 years. Assume Ms. Kremer has made 12 payments and would like to pay off the balance immediately after payment number 12. How much does she need to pay off the remaining balance on her loan?"
Answer #1

Firstly we need to compute the EAR

EAR = (1+ APR/m)^m-1

= (1+4%/365)^365 -1

=0.04080849

Next compute the monthly payments

Using financial calculator

Input:

PV = -(102000-55000) = -47000

N = 12*12 = 144

I/Y = 0.04080849/12

Solve for PMT as 413.35

Next we compute the balance loan

Using financial calculator

Input:

N = 144-12 = 132

I/Y = 0.04080849/12

PMT = - 413.35

Solve for PV as 43,900.30

Balance loan = $43,900.30